# Electron-phonon couplings in locally disordered materials: The case of hybrid halide perovskites

### v1

@@ Abstract
Figure 3(d) shows the one-phonon energy-resolved contribution to the band gap renormalization of MAPbBr3 at $T=300$ K calculated using the Eliashberg function: $g^{2}F_{n}(\omega,T)=\sum_{\nu}\Delta\varepsilon_{n,\nu}(T)\delta(\hbar\omega-\hbar\omega_{\nu}).$ Here, $\hbar\omega_{\nu}$ is the phonon energy of mode $\nu$, $\Delta\varepsilon_{n,\nu}(T)=\frac{1}{2}\frac{\partial^{2}\varepsilon_{n}}{\partial x_{\nu}^{2}}\sigma^{2}_{\nu,T}$ is the mode-resolved energy renormalization of a state $\psi_{n}$ up to second order, and $\sigma^{2}_{\nu,T}$ is the associated mean-square displacement Zacharias and Giustino (2016). $\Delta\varepsilon_{n,\nu}(T)$ is calculated by finite differences using single mode displaced configurations generated by ZG.x Lee et al. (2023). Our results for the reference and polymorphous structures reveal that low energy optical vibrations in the window 2–10 meV and 4–10 meV, respectively, dominate the phonon-induced band gap widening, in agreement with interpretations of previous works Zacharias et al. (2023a); Yazdani et al. (2023); Zhu et al. (2024); Biswas et al. (2024). The modes associated with this energy window have a bending or rocking character, associated with variations in the B-X-B bond angles and octahedral tilting Pérez-Osorio et al. (2015). Higher frequency modes are mostly associated to B-X stretching or molecular librations. Positional polymorphism strongly suppress the positive EPC contribution in the window 2–10 meV, while enhancing the negative contributions in the window 10–20 meV, resulting in significant reduction in the overall gap opening.
To elucidate the origin of this reduction, we investigate the band gap renormalization due to a mode with bending character. If we consider the Kohn-Sham hamiltonian $H_{\rm KS}$, express $\varepsilon_{n}=\braket{\psi_{n}}{H_{\rm KS}}{\psi_{n}}$, take the second derivative with $x_{\nu}$, and apply the Hellmann-Feynman theorem twice Poncé et al. (2014) we obtain:
[…]
Here, $V_{\rm KS}$ is the Kohn-Sham potential and the integration is taken over all electron coordinates ${\bf r}$. To arrive at this expression we also neglected the non-local contribution to $V_{\rm KS}$ Starace (1971). The first and second terms on the right hand side of Eq. (Electron-phonon couplings in locally disordered materials: The case of hybrid halide perovskites) are the Fan-Migdal (FM) and Debye-Waller (DW) EPC coefficients Giustino (2017).
@@
In conclusion, we demonstrate an efficient method for the calculation of temperature-dependent band gaps in complex polymorphous systems based on the ASDM. Our method allows for the simultaneous treatment of anharmonicity and local disorder in both the inorganic and organic sublattices, as well as electron-phonon coupling. We clarify the origin of the gradual band gap blueshift of halide perovskites with temperature, attributing it to the low response of electrons to nuclear vibrations induced by positional polymorphism. Our bottom-up ab initio approach allows for disentangling effects of local disorder, electron-phonon coupling, and thermal expansion, enabling our understanding of the individual role of these key phenomena in optoelectronic properties. In addition, we elucidate the key role of the A-site cation in influencing local disorder, band gap, and the electron-phonon renormalization. In the future, it will be interesting to investigate temperature-dependent properties of layered halide perovskites and heterostructures Sidhik et al. (2022); Metcalf et al. (2023); Jiang et al. (2025). Our approach also forms the basis for anharmonic electron-phonon calculations in polymorphous solids, including the study of phonon-assisted optical absorption Zacharias et al. (2015); Noffsinger et al. (2012); Zacharias and Giustino (2016), carrier mobilities Herz (2017); Schlipf et al. (2018); Poncé et al. (2019); Cucco et al. (2024), polarons Franchini et al. (2021); René de Cotret et al. (2022); Lafuente-Bartolome et al. (2022); Zhang et al. (2023); Seiler et al. (2023); Biswas et al. (2024), and exciton-phonon interactions Alvertis et al. (2023); Dai et al. (2024b).
@@ Acknowledgements.
This research was funded by the European Union (project ULTRA-2DPK / HORIZON-MSCA-2022-PF-01 / Grant Agreement No. 101106654). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. We thank D. R. Ceratti for graciously providing data on temperature-dependent band gaps of Br-based hybrid halide perovskite single crystals. J.E. acknowledges financial support from the Institut Universitaire de France. F.G. was supported by the the Robert A. Welch Foundation under Award No. F-2139-20230405 and by the National Science Foundation under DMREF Grant No. 2119555. G.V. acknowledges funding from the ANR through the CPJ program and the SURFIN project (ANR-23-CE09-0001), the ALSATIAN project (ANR-23-CE50-0030), and the ANR under the France 2030 programme, MINOTAURE project (ANR-22-PETA-0015). We acknowledge that the results of this research have been achieved using computational resources from the EuroHPC Joint Undertaking and supercomputer LUMI [https://lumi-supercomputer.eu/], hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Extreme Scale Access call.

### v2

@@ Abstract
Figure 3(d) shows the one-phonon energy-resolved contribution to the band gap renormalization of MAPbBr3 at $T=300$ K calculated using the Eliashberg function: $g^{2}F_{n}(\omega,T)=\sum_{\nu}\Delta\varepsilon_{n,\nu}(T)\delta(\hbar\omega-\hbar\omega_{\nu}).$ Here, $\hbar\omega_{\nu}$ is the phonon energy of mode $\nu$, $\Delta\varepsilon_{n,\nu}(T)=\frac{1}{2}\frac{\partial^{2}\varepsilon_{n}}{\partial x_{\nu}^{2}}\sigma^{2}_{\nu,T}$ is the mode-resolved energy renormalization of a state $\psi_{n}$ up to second order, and $\sigma^{2}_{\nu,T}$ is the associated mean-square displacement Zacharias and Giustino (2016). $\Delta\varepsilon_{n,\nu}(T)$ is calculated by finite differences using single mode displaced configurations generated by ZG.x Lee et al. (2023). Our results for the reference and polymorphous structures reveal that low energy optical vibrations in the window 2–10 meV and 4–10 meV, respectively, dominate the phonon-induced band gap widening, in agreement with interpretations of previous works Zacharias et al. (2023a); Yazdani et al. (2023); Zhu et al. (2024); Biswas et al. (2024). The modes associated with this energy window have a bending or rocking character associated with octahedral tilting Pérez-Osorio et al. (2015). Higher frequency modes are mostly associated to B-X stretching or molecular librations. Positional polymorphism strongly suppress the positive EPC contribution in the window 2–10 meV, while enhancing the negative contributions in the window 10–20 meV, resulting in significant reduction in the overall gap opening.
To elucidate the origin of this reduction, we investigate the band gap renormalization due to a mode with bending character. If we consider the Kohn-Sham Hamiltonian $H_{\rm KS}$, express $\varepsilon_{n}=\braket{\psi_{n}|H_{\rm KS}|\psi_{n}}$, take the second derivative with $x_{\nu}$, and apply the Hellmann-Feynman theorem twice Poncé et al. (2014) we obtain:
$\hskip-11.38092pt\frac{\partial^{2}\varepsilon_{n}}{\partial x_{\nu}^{2}}=\int d{\bf r}\frac{\partial V_{\rm KS}}{\partial x_{\nu}}\frac{\partial|\psi_{n}({\bf r})|^{2}}{\partial x_{\nu}}+\int d{\bf r}\frac{\partial^{2}V_{\rm KS}}{\partial x_{\nu}^{2}}|\psi_{n}({\bf r})|^{2}.$ (2)
Here, $V_{\rm KS}$ is the Kohn-Sham potential and the integration is taken over all electron coordinates $\bf r$}. To arrive at this expression we also neglected the non-local contribution to $V_{\rm KS}$ Starace (1971). The first and second terms on the right hand side of Eq. (2) are the Fan-Migdal (FM) and Debye-Waller (DW) EPC coefficients Giustino (2017).
@@
In conclusion, we generalize the ASDM for the simultaneous treatment of anharmonicity and configurational entropy in both the inorganic and organic sublattices, as well as electron-phonon coupling, without long path-integral or classical MD runs. We clarify the origin of the gradual band gap blueshift of halide perovskites with temperature, attributing it to the low response of electrons to nuclear vibrations induced by positional polymorphism. We elucidate the influence of the A-site cation on local disorder, band gap, and electron-phonon renormalization, establishing rules for materials engineering. In the future, it will be interesting to investigate temperature-dependent properties of emerging hybrid materials with optoelectronic, ferroelectric, and thermoelectric applications Haque et al. (2020); Sidhik et al. (2022); Zhang et al. (2022); Dirin et al. (2023); Metcalf et al. (2023); Jiang et al. (2025). Our approach forms the basis for electron-phonon calculations in polymorphous solids, including phonon-assisted absorption Zacharias et al. (2015); Noffsinger et al. (2012); Zacharias and Giustino (2016), mobilities Herz (2017); Schlipf et al. (2018); Poncé et al. (2019); Cucco et al. (2024), polarons Franchini et al. (2021); René de Cotret et al. (2022); Lafuente-Bartolome et al. (2022); Zhang et al. (2023); Seiler et al. (2023); Biswas et al. (2024), and exciton-phonon interactions Alvertis et al. (2023); Dai et al. (2024b).
Data leading to the results of this study is available in the NOMAD repository Zacharias (2025).
@@ Acknowledgements.
This research was funded by the European Union (project ULTRA-2DPK / HORIZON-MSCA-2022-PF-01 / Grant Agreement No. 101106654). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. We thank D. R. Ceratti for graciously providing data on temperature-dependent band gaps of Br-based hybrid halide perovskite single crystals. J.E. acknowledges financial support from the Institut Universitaire de France. F.G. was supported by the Robert A. Welch Foundation under Award No. F-2139-20230405 and by the National Science Foundation under DMREF Grant No. 2119555. G.V. acknowledges funding from the ANR through the CPJ program and the SURFIN project (ANR-23-CE09-0001), the ALSATIAN project (ANR-23-CE50-0030), and the ANR under the France 2030 programme, MINOTAURE project (ANR-22-PETA-0015). We acknowledge that the results of this research have been achieved using computational resources from the EuroHPC Joint Undertaking and supercomputer LUMI [https://lumi-supercomputer.eu/], hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Extreme Scale Access call.